OpenStudy (anonymous):

The circumference of a circle is 15. Use the formula bellow to find the radius r. Use 3.14 as an approximation for 3.14 and round your answer to the nearest hundredth.

OpenStudy (anonymous):

that's the formula

OpenStudy (aravindg):

C= 2 pi r Divide both sides by 2pi to get r

OpenStudy (anonymous):

r = 5.21 ?

OpenStudy (aravindg):

Try again?

OpenStudy (anonymous):

alright i think i did something wrong

OpenStudy (anonymous):

can you help me wiht this im not sure how to break this down

OpenStudy (aravindg):

Just put the values C=15 2 pi=2 *3.14

OpenStudy (aravindg):

C/2pi=?

OpenStudy (anonymous):

We Stupid Ok xc

OpenStudy (anonymous):

4.78?

OpenStudy (aravindg):

Nope.

OpenStudy (aravindg):

2*pi=?

OpenStudy (anonymous):

6.28

OpenStudy (aravindg):

15/6.28=?

OpenStudy (anonymous):

2.39 :D

OpenStudy (anonymous):

thanks buddy :)

OpenStudy (aravindg):

yw :)